OpenStudy (anonymous):

Look at the table. Make a conjecture about the sum of the first 20 positive even numbers. The sum is 21 × 22. The sum is 20 × 21. The sum is 20 × 20. The sum is 19 × 20. can someone help me lol

OpenStudy (anonymous):

\[2+4=6\\ 2+4+6=12\\ 2+4+6+8=20\\ 2+4+6+8+10=30\]